Dot or Scalar Product

1. Calculate the dot product of vectors a and b where a is 4 units and b is 8 units and angle Theta is 67
1
Expert's answer
2021-10-06T15:53:26-0400

By definition, the dot product is:


"(\mathbf{a},\mathbf{b}) = ab\cos\theta"

where "a = 4, b = 8" are the magnitudes of the vectors, and "\theta = 67\degree" is the angle between them. Thus, obtain:


"(\mathbf{a},\mathbf{b}) = 4\cdot 8\cdot \cos67\degree \approx 12.5"

Answer. 12.5.